OpenStudy (anonymous):

WIll give medal and fan A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

I need help understanding this question and the steps to complete it xD

OpenStudy (kira_yamato):

|x − 0.525| = 0.01 => x - 0.525 = -0.01 OR x - 0.525 = 0.01 (by definition of modulus) Therefore x = -0.01 + 0.525 = 0.515 or x = 0.01 + 0.525 = 0.535 In other words, diameter is in the interval with min 0.515 and max 0.535

OpenStudy (anonymous):

thanks for the help